OpenStudy (anonymous):

given ABC=PQR, m B= 4v+5, and m Q= 5v-6, find m b and m q

OpenStudy (alexandervonhumboldt2):

4v+5=5v-6 -v=-11 v=11

OpenStudy (anonymous):

then do i plug in 11 for both sides

OpenStudy (alexandervonhumboldt2):

4*11+5=5*11-6

OpenStudy (alexandervonhumboldt2):

49=49!!!

OpenStudy (anonymous):

thx